OpenStudy (anonymous):

Graph the following line by plotting the x and y intercepts: 6x – 4y = -8

OpenStudy (anonymous):

your first step is to convert this into y = mx +b form.

OpenStudy (anonymous):

do that by first subtracting the 6x on both sides so that your equation is now -4y = -6x -8

OpenStudy (anonymous):

and, word of advice here, i would multiply both sides of your equation by negative 1 now so that you have 4y = 6x + 8. just because positive numbers are generally easier to solve with

OpenStudy (anonymous):

so how would i graph it then?

OpenStudy (anonymous):

now divide both sides of your equation by 4 so that you make the y be on its own. your equation should by y = 3/2x + 2

OpenStudy (anonymous):

^^ is in y = mx + b form. when in this form, your b is your y - intercept

OpenStudy (anonymous):

so your y is (0, 2)

OpenStudy (anonymous):

ok so are there any negatives in it?

OpenStudy (anonymous):

yes, the x will be negative and i will show you why

OpenStudy (anonymous):

(-1,0) right

OpenStudy (anonymous):

ok got it thx

OpenStudy (anonymous):

sorry, it isnt (-1, 0)

OpenStudy (anonymous):

(-1,0) is x

OpenStudy (anonymous):

no, sorry because your next point is (-2, -1)

OpenStudy (anonymous):

down 3, over 2

OpenStudy (anonymous):

i got it correct y=(0,2) and x=(-1,0)

OpenStudy (anonymous):

yah down 3 which takes you to 0 the over to left 1

OpenStudy (anonymous):

i got something different, i ended up getting -4/3 as my x

OpenStudy (anonymous):

which is -1

OpenStudy (anonymous):

oh well it was -1,0 XD thx for the help man

OpenStudy (anonymous):

sure thing